OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (anonymous):

if two lines are parallel, then both lines yield the same slope.

OpenStudy (anonymous):

so our new line would be in the form y=2x+c simply input the point given (2,-1) into that new equation to find c

OpenStudy (anonymous):

i think D

OpenStudy (anonymous):

-1=2(2)+c c=-1-4 y=2x-5

OpenStudy (anonymous):

thank you for helping :)

OpenStudy (anonymous):

You know how the equation has x and y in it. and you know how (in this case) x = 2 and y = -1, well. you want to plug those numbers in the x and y slots. You can do that for each, or as chris00 said, you can plug it in that equation to find c.

OpenStudy (anonymous):

thank you makes sense now

OpenStudy (anonymous):

No problem. :)